# How typical is contextuality?

### v1

@@ 2.3 Comparison to typicality of Bell nonlocality
Contextuality is closely related to a type nonclassicality, called nonlocality bell1964on ; Brunner2014 . Nonlocality arises in Bell scenarios, where statistics generated by spacelike-separated measurements on systems prepared in entangled states may sometimes resist a classical explanation. In Bell scenarios, the outcome statistics are usually referred to as correlations. For two parties (Alice and Bob) in a Bell scenario, the quantum correlations arise via Born’s rule: $p(ab|xy)=\mathrm{tr}\left\{(M_{a|x}\otimes M_{b|y})\rho_{AB}\right\}$. Here, $M_{a|x}$ (resp. $M_{b|y}$) are the POVM elements of Alice’s (resp. Bob’s), and $\rho_{AB}$ is the density matrix representing the quantum state of the shared system. Previous work has investigated whether nonlocality is a typical property of pure entangled states deRosier2017 ; de2020strength ; Lipinska2018 . The general conclusion is that as either the number of parties or the number of measurement settings increases, typicality of nonlocality exceeds $99\%$ for pure states and projective measurements. Moreover, it has been shown that for any pure bipartite entangled state, typicality tends to unity as the number of measurement settings tends to infinity Lipinska2018 .
@@ 3 Application to parity-oblivious multiplexing
Case ii) The decoding party does not implement random projective measurements but rather random POVMs. This may happen when the decoding party also lacks any control over the sharpness of the measurements or when the channel between the encoding and the decoding parties is subject to noise. The typicality of contextuality in this case is equal to $55.5\%$, which is significantly lower than for case i) with random projective measurements. The results of the analysis of the average success rate are summarized below.
@@
Consider the prepare-and-measure scenario $(n=8,m=3,d=2;N=10^{6})$, with $m=3$ randomly sampled dichotomic POVMs but the $n=8$ fixed states used in the optimal strategy for POM. The typicality of contextuality for this scenario is $t_{(n,m,d;N)}\approx 55.5\%$. The average success rate in the POM task is $\bar{s}=0.68$ ($\sigma_{s}=0.02$), which is about $2\%$ above the optimal noncontextual strategy. The average robustness of contextuality is $\bar{r}=0.06$ ($\sigma_{r}=0.07$).
Unsurprisingly, the strategy with random POVMs performs very poorly. Although the success rate is higher than the classical one in over half of the cases, it is only negligibly higher on average, with only a $2\%$ improvement over the noncontextual strategy.
@@
Random POVMs 55.5% 0.06 2%
Table 1: Typicality of contextuality, average advantage over the optimal classical strategy, and average robustness of contextuality to depolarizing noise of 3-to-1 parity-oblivious multiplexing tasks in a qubit with optimal encoding strategy but suboptimal measurements. The first row includes the data in the optimal quantum scenario, for reference. Average advantages had a variance of about $\sigma_{s}=0.02$ for all scenarios, and average robustness, of about $\sigma_{r}=0.07$.
@@ Acknowledgments
We thank Paweł Cieśliński for discussions about typicality of Bell nonlocality and Piotr Mironowicz for the support with the cluster. V.P.R., B.Z., and R.D.B acknowledge support by the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, GA No. 101070558, funded by the European Union, NSERC (Canada), and UKRI (UK). D.S. and R.D.B. were supported by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. JHS was funded by the European Commission by the QuantERA project ResourceQ under the grant agreement UMO2023/05/Y/ST2/00143. This work is partially carried out under IRA Programme, project no. FENG.02.01-IP.05-0006/23, financed by the FENG program 2021-2027, Priority FENG.02, Measure FENG.02.01., with the support of the FNP. We acknowledge the use of a computational server financed by the Foundation for Polish Science (IRAP project, ICTQT, contract no. 2018/MAB/5/AS-1, co-financed by EU within Smart Growth Operational Programme). Some figures were prepared using Mathcha.
@@ B.1 Random pure states and unitaries
which specifies how the angles must be sampled with the correct weight in order to obtain Haar-random unitaries Zyczkowski1994 .
B.2 Random mixed states
@@
It is worth mentioning that when we sample random POVM elements, we do so by sampling random density matrices as discussed above. This does not sample uniformly over the set of POVM elements, since we are restricting ourselves to normalized ones. Although this does not impact the assessments of contextuality provided by the linear program selby2024linear (as nonclassicality depends only on the cone generated by such effects selby23fragments ), an interesting question is how to design a more meaningful sampling method for effects. The questions of how such a method is designed and implemented and what its impact is on typicality assessments are deferred to a future version of this work.
@@ Appendix C Technical discussion
Appendix D Open-source repository
@@
The repository requires qutip version 5.2 or above. It samples random states as described in Sec. 1.2 and Appendix B through the functions random_density_matrix, random_effects, fixed_effects and random_unitary. In particular, random_density_matrix will sample either (i) a pure state uniformly via Haar measure if pure = True is provided as an input or (ii) a full-rank mixed density operator by Ginibre sampling, employing the exclusion criterion if purity lies out of the interval determined by the inputs (upperbound and lowerbound). A new version of this script is in progress and will provide additional tools to sample POVMs with unnormalized elements and with multiple outcomes.

### v2

@@ 2.3 Comparison to typicality of Bell nonlocality
Contextuality is closely related to a type nonclassicality, called nonlocality bell1964on ; Brunner2014 . Nonlocality arises in Bell scenarios, where statistics generated by spacelike-separated measurements on systems prepared in entangled states may sometimes resist a classical explanation. In Bell scenarios, the outcome statistics are usually referred to as correlations. For two parties (Alice and Bob) in a Bell scenario, the quantum correlations arise via Born’s rule: $p(ab|xy)=\mathrm{tr}\left\{(M_{a|x}\otimes M_{b|y})\rho_{AB}\right\}$. Here, $M_{a|x}$ (resp. $M_{b|y}$) are the POVM elements of Alice’s (resp. Bob’s), and $\rho_{AB}$ is the density matrix representing the quantum state of the shared system. Previous work has investigated whether nonlocality is a typical property of pure entangled states Drumond2012 ; deRosier2017 ; Lipinska2018 ; Drumond2018 ; de2020strength . The general conclusion is that as either the number of parties or the number of measurement settings increases, typicality of nonlocality exceeds $99\%$ for pure states and projective measurements. Moreover, it has been shown that for any pure bipartite entangled state, typicality tends to unity as the number of measurement settings tends to infinity Lipinska2018 , while actual violations of Bell inequalities grow increasingly small with the local dimension of the quantum system Drumond2012 ; Drumond2018 .
@@ 3 Application to parity-oblivious multiplexing
Case ii) The decoding party does not implement random projective measurements but rather random POVMs. This may happen when the decoding party also lacks any control over the sharpness of the measurements or when the channel between the encoding and the decoding parties is subject to noise. The typicality of contextuality in this case is equal to $15.2\%$, which is significantly lower than for case i) with random projective measurements. The results of the analysis of the average success rate are summarized below.
@@
Consider the prepare-and-measure scenario $(n=8,m=3,d=2;N=10^{6})$, with $m=3$ randomly sampled dichotomic POVMs but the $n=8$ fixed states used in the optimal strategy for POM. The typicality of contextuality for this scenario is $t_{(n,m,d;N)}\approx 15.2\%$. The average success rate in the POM task is $\bar{s}=0.669$ ($\sigma_{s}=0.006$), which is about $0.3\%$ above the optimal noncontextual strategy. The average robustness of contextuality is $\bar{r}=0.01$ ($\sigma_{r}=0.03$).
@@
Random POVMs 15.2% 0.01 0.3%
Table 1: Typicality of contextuality, average advantage over the optimal classical strategy, and average robustness of contextuality to depolarizing noise of 3-to-1 parity-oblivious multiplexing tasks in a qubit with optimal encoding strategy but suboptimal measurements. The first row includes the data in the optimal quantum scenario, for reference. Average advantages had a variance of about $\sigma_{s}=0.02$ for the scenarios with rotated optimal measurements and with random projective measurements, while $\sigma_{s}=0.006$ for the scenario with random POVMs. For the average robustness, the variance was of about $\sigma_{r}=0.07$ for the former cases, and $\sigma_{r}=0.03$ for the latter.
@@ Acknowledgments
We thank Paweł Cieśliński for the discussions on typicality of Bell nonlocality and Piotr Mironowicz for the support with the cluster. V.P.R., B.Z., and R.D.B acknowledge support by the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, GA No. 101070558, funded by the European Union, NSERC (Canada), and UKRI (UK). D.S. and R.D.B. were supported by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. JHS was funded by the European Commission by the QuantERA project ResourceQ under the grant agreement UMO2023/05/Y/ST2/00143. This work is partially carried out under IRA Programme, project no. FENG.02.01-IP.05-0006/23, financed by the FENG program 2021-2027, Priority FENG.02, Measure FENG.02.01., with the support of the FNP. We acknowledge the use of a computational server financed by the Foundation for Polish Science (IRAP project, ICTQT, contract no. 2018/MAB/5/AS-1, co-financed by EU within Smart Growth Operational Programme). Some figures were prepared using Mathcha.
@@ B.1 Random pure states and unitaries
[…]
B.2 Random mixed states and POVMs
@@
In order to sample dichotomic POVMs, we employ the method defined in Ref. kukulski2021 . That is, we sample pairs $(G_{1},G_{2})$ of $d\times d$ complex matrices belonging to the Ginibre ensamble, and then proceed to construct the positive semidefinite elements
$E_{1}=G_{1}G_{1}^{\dagger},\quad E_{2}=G_{2}G_{2}^{\dagger},\quad S=E_{1}+E_{2}.$ (23)
The POVM is therefore given by the normalized elements
$M=\{M_{1}:=S^{-1/2}E_{1}S^{-1/2},\ M_{2}:=S^{-1/2}E_{2}S^{-1/2}\},$ (24)
which ensures that the elements sum up to the identity. As demonstrated in Ref. kukulski2021 , this measure can be generalised to quantum channels and is equivalent to other commonly considered measures. The implementation we provide in Ref. github also admits sampling POVMs with a larger number of outcomes.
To sample POVMs with bounded sharpness, we employ the same rejection method as the one employed for density operators, with the caveat that for POVMs the quantity to be verified is given by the sharpness $\eta$ Salehi2021 , i.e., for a POVM $M:=\{M_{i}\}_{i=1}^{k}$ in a $d$-dimensional quantum system,
$\eta_{M}:=\frac{1}{d}\sum_{i=1}^{k}\mathrm{tr}\left\{M_{i}^{2}\right\}.$ (25)
@@ Appendix C Technical discussion
Appendix D Analytical proofs
We begin by proving Proposition 2.
See 2
Proof.
Notice that for $n\leq d^{2}$ or $m\leq d^{2}/2$ (in the case of dichotomic measurements), $t_{(n,m,d)}=0$ by Lemma 1, so we now demonstrate that $t_{(n,m,d)}<1$ for the remaining cases.
Consider the GPT described by the full set of quantum states $\Omega$ for a quantum system of dimension $d$, and its dual $\mathcal{E}$ representing the full set of effects. Hereon, $\Omega$ will be treated as a compact convex body in its affine state space of real dimension $d^{2}-1$. Fix a sampling model that uniformly samples states in $\Omega$ and effects in $\mathcal{E}$, then for a randomly sampled finite set of effects $\mathcal{E}_{m}\subsetneq\mathcal{E}$ with $m>d^{2}/2$ the dual polytope $\mathcal{E}_{m}^{*}$ strictly contains $\Omega$ and is full-dimensional with nonzero probability.
Since $\mathcal{E}_{m}^{*}$ generically strictly contains $\Omega$, it is possible to select $v_{0}\in\mathcal{E}_{m}^{*}\backslash\Omega$. Moreover, it is always possible to choose a full-dimensional simplex $\triangle\subsetneq\mathcal{E}_{m}^{*}$ containing $v_{0}$ as one of its vertices and the remaining vertices in $\Omega$, such that $\mathsf{Int}(\triangle)\cap\mathsf{Int}(\Omega)\neq\emptyset$. By construction, $\mathsf{Int}(\triangle)\cap\mathsf{Int}(\Omega)\subsetneq\triangle$ and $\mathcal{E}_{m}\subsetneq\triangle^{*}$, so any GPT fragment contained in these sets of states and effects will be simplex-embeddable.
By construction, there exists a boundary point $\omega\in\partial\Omega$ such that $\omega\in\mathsf{Int}(\triangle)$. Therefore, there is an $\varepsilon$-ball $B_{\varepsilon}(\omega)\subsetneq\triangle$ around $\omega$. It follows that, given $\mathsf{Vol}$ the Lebesgue volume in $\mathbb{R}^{d^{2}-1}$, $\Omega_{\varepsilon}:=B_{\varepsilon}(\omega)\cap\Omega$ contains interior points of $\Omega$ and thus will have positive Lebesgue volume. Therefore, there is a finite probability of sampling a state in this region given by
$p:=\frac{\mathsf{Vol}[\Omega_{\varepsilon}]}{\mathsf{Vol}[\Omega]}>0,$ (27)
and, by independence, the probability of sampling $\Omega_{n}$ quantum states in this region is also positive and given by $p^{n}$. With probability $p^{n}>0$, then, we end up sampling a set of pure quantum states that satisfies $\Omega_{n}\subset\triangle$, and therefore corresponds to a simplex-embeddable fragment. The typicality of contextuality is therefore bounded by
$t_{(n,m,d)}\leq 1-p^{n}<1.$ (28)
[…]
By construction, there is a boundary point $\omega\in\partial\Omega$ such that $\omega\in\mathsf{Int}(\triangle)$. Therefore, there is an $\varepsilon$-ball $B_{\varepsilon}(\omega)\subset\triangle$ around $\omega$. It follows that, given $\mu$ the Haar measure in $\partial\Omega$, $\Omega_{\varepsilon}:=B_{\varepsilon}(\omega)\cap\partial\Omega$ contains points of $\partial\Omega$ and thus will have positive Haar measure. Therefore, there is a finite probability of sampling a state in this region, given by
$p:=\frac{\mu(\Omega_{\varepsilon})}{\mu(\Omega)}>0,$ (29)
and by independence, the probability of sampling $\Omega_{n}$ pure states in this region is also positive and given by $p^{n}>0$, so the bound on typicality takes on the same form as Eq. 28. ∎
Naturally, the upper bound in Eq. (28) is not tight: there are infinitely many such simplices, and in this proof we restricted the scope of simplex-embeddings to those where the simplices have the same dimension as the sampled set of quantum states.
We now proceed to prove Proposition 3.
See 3
Proof.
Let us begin by proving Eq. (7).
Similarly to the previous proof, consider the GPT described by the full set of quantum states $\Omega$ for a quantum system of dimension $d$, and its dual $\mathcal{E}$ representing the full set of effects. Hereon, $\Omega$ and $\mathcal{E}$ will be treated as a compact convex bodies in their affine spaces of real dimension $d^{2}-1$. Notice that for any $d$, there is no simplex $\triangle$ and embeddings $\iota:\Omega\to\triangle$ and $\kappa:\mathcal{E}\to\triangle$ such that $(\iota(\Omega),\kappa(\mathcal{E})\subseteq(\triangle,\triangle^{*})$.
Fix a sampling model that uniformly samples pure states in $\partial\Omega$ and pure effects in $\partial\mathcal{E}$. From Ref. zhang2025reassessing , any set $\Omega_{n}$ of $n>d^{2}$ randomly sampled states will be such that the fragment $(\Omega_{n},\mathcal{E})$ does not admit of a simplex embedding. Assume that there is $\triangle$ and $\iota,\kappa$ such that the fragment $(\Omega_{n},\mathcal{E}^{\prime})$ admits of a simplex embedding for some convex hull $\mathcal{E}^{\prime}$ of pure effects. Then the result in Ref. zhang2025reassessing imposes that $\mathcal{E}^{\prime}\neq\mathcal{E}$, and since $\mathcal{E}$ is the set of all possible effects for this system, it must be that $\mathcal{E}^{\prime}\subsetneq\mathcal{E}$, and therefore $\partial\mathcal{E}^{\prime}\subsetneq\partial\mathcal{E}$. This means that $\partial\mathcal{E}^{\prime}\cap\partial\mathcal{E}=\partial$$\mathcal{E}^{\prime}$. It follows that, given $\mu$ the Haar measure in $\partial\mathcal{E}$, the likelihood of sampling an effect in $\partial\mathcal{E}^{\prime}$ is
$p_{\Omega_{n},\triangle}:=\frac{\mu(\partial\mathcal{E}^{\prime})}{\mu(\partial\mathcal{E})}<1.$ (30)
The probability of finding such an effect for any given $\Omega_{n}$ will be given by maximising $p_{(\Omega_{n},\triangle)}$ over all possible simplex-embeddings, and then averaging that over all possible $\Omega_{n}$. That is,
$p:=\int d\mu(\Omega_{n})\max_{\triangle,\iota,\kappa}\frac{\mu(\partial\mathcal{E}^{\prime})}{\mu(\partial\mathcal{E})}<1.$ (31)
The probability of independently sampling $m$ effects not in this setup is therefore given by
$t_{(n,m,d)}=1-p^{m},$ (32)
and in the asymptotic limit,
$\lim_{m\to\infty}t_{(n,m,d)}=1.$ (33)
The proof follows similarly for Eq. 8. From the result in Ref. zhang2025reassessing , for any set $\mathcal{E}_{m}$ of dichotomic pure effects with $m\geq d^{2}/2$, linear dependences will be found among them and so $(\Omega,\mathcal{E}_{m})$ will not admit of a simplex embedding. Assume therefore that there is a $\triangle$ and $\iota,\kappa$ such that $(\Omega^{\prime},\mathcal{E}_{m})$ admits of a simplex embedding for some convex hull $\Omega^{\prime}$ of pure states. The result in Ref. zhang2025reassessing imposes that $\Omega^{\prime}\neq\Omega$, and by the same reasoning as previously it follows that $\partial\Omega^{\prime}\cap\partial\Omega=\partial\Omega^{\prime}$, so the likelihood of sampling a state in $\partial\Omega^{\prime}$ given the Haar measure $\mu$ in $\partial\Omega$ is
$p_{\mathcal{E}_{m},\triangle}:=\frac{\mu(\partial\Omega^{\prime})}{\mu(\partial\Omega)}<1.$ (34)
Again we must maximise it over all possible simplex embeddings and then average over all possible sets $\mathcal{E}_{m}$, but this probability remains lower than one:
$p:=\int d\mu(\mathcal{E}_{m})\max_{\triangle,\iota,\kappa}\frac{\mu(\partial\Omega^{\prime})}{\mu(\partial\Omega)}<1,$ (35)
so the probability of independently sampling $n$ states not in this setup is therefore given by
$t_{(n,m,d)}:=1-p^{n},$ (36)
so in the asymptotic limit
$\lim_{n\to\infty}t_{(n,m,d)}=1.$ (37)
Appendix E Open-source repository
@@
The repository requires qutip version 5.2 or above. It samples random states as described in Sec. 1.2 and Appendix B through the functions random_density_matrix, random_effects, fixed_effects and random_unitary. In particular, random_density_matrix will sample either (i) a pure state uniformly via Haar measure if pure = True is provided as an input or (ii) a full-rank mixed density operator by Ginibre sampling, employing the exclusion criterion if purity lies out of the interval determined by the inputs (upperbound and lowerbound). Similarly, random_effects will sample either (i) $m$ projective measurements via Haar measure if pure = True; or (ii) $m$ POVMs with $k$ outputs (with $k=2$ by default) by Ginibre sampling, employing the exclusion criterion if the sharpness (Eq. 25) lies out of the interval determined by the inputs (upperbound and lowerbound).